OpenStudy (anonymous):

Help please!!! (: A net force of 2.6N(i-hat) - 2.5N(j-hat) acts on a 0.9 kg object. Find the acceleration a. =>_______m/s^2(i-hat) + ________m/s^2(j-hat)

OpenStudy (anonymous):

\[F - F \mu = ma\] \[2.6N - 2.1N = 0.9kg(a)\] \[0.5 = 0.9a\] \[a = \frac{ 0.5 }{ 0.9 }\] \[a = 0.556 m/s ^{2}\]

OpenStudy (loser66):

you have \[\vec F = m\vec a\\\vec a=\frac{\vec F}{m}=\frac{(2.6 \hat{i} -2.5\hat{j}}{0.9}= 2.9\hat{i}-2.8\hat{j} (m/s^2)\] that means to horizontal direction, the acceleration is 2.9 m/s^2 and to the vertical direction, the acceleration is 2.8 m/s^2